# Resolvi: A Reference Architecture for Scalable and Interoperable Entity Resolution

### v2

@@ Abstract
Entity resolution (ER) is a critical task in data management which identifies whether multiple records refer to the same real-world entity. Despite its significance across domains such as healthcare, finance, and machine learning, implementing effective ER systems remains challenging due to the abundance of methodologies and tools, leading to a paradox of choice for practitioners. This paper proposes Resolvi, a reference architecture which prioritizes extensibility, interoperability, and scalability in ER systems. By analyzing existing ER frameworks and literature, we establish a structured approach to designing ER solutions that address common challenges. Additionally, we explore best practices for system implementation and deployment strategies to facilitate large-scale entity resolution. Through this work, we aim to provide a foundational blueprint that assists researchers and practitioners in developing robust, scalable ER systems while reducing the complexity of architectural decisions.
journal: Systems and Software
@@ 2.1 Related Work
In the absence of a comprehensive review focusing on the disambiguation aspect of named entity recognition and disambiguation, several independent sources were studied to better understand the problem. The opinion that named entity disambiguation systems are designed very similarly to merge/purge systems is rooted in the NERD framework [56] itself, which implements a special ontology-based merge algorithm of extracted named entities. FOX [64] uses building blocks which are prevalent in record linkage systems, too (e.g indexing components, external knowledge bases). Modern systems such as Ensemble Nerd [9] use separate black box subsystems (in this case neural networks) to perform feature extraction and matching, similarly to Ditto [35]. Finally, to the best of our knowledge there is no dispute over the need to extract structured information from unstructured text across the named entity recognition literature.
@@ 2.4 Reference ER Systems
Most of the extensibility built into existing systems focuses on the key aspects of ER (matching, clustering, data preprocessing). The Magellan [30] framework for entity matching is extensible in a completely different way: it is allows using matching components such as Ditto [35] or replacing the entire rule-based ER runtime with a full neural network architecture designed for ER such as DeepMatcher [46]. This emphasizes the higher extensibility of ER systems with a framework design.
@@ 5.4 Matching Engine
Additionally, the matching process can be symbolically reduced to a functional representation, barring the need to inspect the detailed logical components involved. The black box approach to matching starting with the Stanford Entity Resolution Framework (SERF) [3]. One critique of the SERF match function is that it implicitly assumes a fixed output structure, whereas in practice, matching strategies differ widely.
@@
System architectures should design the matching engine in the most suitable way to their problem domain and chosen ER model. Certain architecture may benefit from leveraging multiple matching engines. A legitimate concern may regard the ability to scale the matching engine independently of the other components in the ER pipeline because of the higher computational expense inherent in this component. Finally, the possibility to implement systems that do not leverage similarity, renders the matching engine as an optional component in the ER pipeline. These considerations render the Strategy design pattern [19] useful for encapsulating the algorithms which sit at the core of the matching engine.
5.5 Reference Store
The Reference Store appears in both incremental and batch ER, even though it is strictly necessary only for the incremental variant. Existing systems often implement it as an in-memory structure. Common representations include a sequence of pairs of matching entity references [35], a similarity graph [53] or a logical graph [60]. Certain closed-source implementations reportedly persist long-lived identifiers linked to entity references [70], though their exact design remains unknown.
When it is designed as a component, it is reminiscent of the Memento design pattern [19] because it is primarily used to store and access previous versions of the same data. Common concerns to consider when designing the reference store include data availability and integrity.
@@ 5.6 Clustering Engine
Widely regarded as an essential step, not all reference systems implement it transparently. Some ER systems, such as Duke [21] obfuscate clustering by making it part of intermediate storage operations which use its system of MatcherListener. System developers are informed about clustering possibilities only once the LinkDatabaseMatchListener and LinkDatabase concepts are perused. This is the only example of obfuscating the pipeline structure of the ER process among the reference ER systems. On the other hand, some entity matching systems such as Ditto [35] or DeepMatcher [46] focus solely on identifying matching entity references, partially or completely omitting transitive resolution. In some cases, this approach can be beneficial, as it offloads the bad-triplet problem [1] to downstream systems.
At the other end of the spectrum, entity alignment and record linkage systems designed around graph models inherently support clustering. Some ER solutions go further by eliminating the matching phase entirely. For example, clustering subsumes matching by employing functional similarity measures and a carefully chosen merge function [3], generic distance metrics [71], or probabilistic models which form clusters directly from the comparison space using similarity functions to model colocation (D-blink [40]).
When applied, clustering typically operates on data supplied by the reference store. In cases where matching is omitted, the entity reference store effectively defines the relevant comparison space. Systems such as the JedAI [53] and FAMER [60] already support various clustering algorithms [50, 61], offering flexible options for different ER workflows.
Given its varying adoption across reference systems and the limited research on cluster coherence [70], system architectures may design a clustering engine if their specific problem domain and ER model of choice require one. In object-oriented designs, the Strategy design pattern [19] may be used to encapsulate the various algorithms sitting at the core of the clustering engine.
New system architectures should carefully consider clustering engines that work with incremental ER. Only some clustering algorithms (such as agglomerative clustering) are generally incremental, i.e can be easily carried out using the batch linkage method [24].
Finally, the evaluation of clustering performance is not as common among the reference ER systems as the fundamental importance of the clustering subsystem to ER would imply. D-blink [40] is an example system which implements this feature. ER system architectures which incorporate clustering engines should also provide the means for evaluating clustering performance along with matching performance. Future efforts in this area would improve ER system transparency and reliability.
@@ 6 ER Runtime Process
Up until this point, the paper focused on generic use cases and abstract logical components. Procedurally ER may be interpreted as a series of subprocesses and activities which, in abstract, remain consistent across existing approaches.
@@ 6.2 Comparison Space Generation
The process of generating the comparison space, denoted as (C) in Figure 10, consists of three main phases: candidate arrangement, comparison input generation, and filtering. Candidate arrangement reorganizes entity references into efficient data structures without loss of information. Techniques such as blocking, indexing, and graph edge modification exemplify this step, resulting in structures like block tables, balanced trees, or graphs. In some cases, these structures may be persistently stored for later retrieval.
@@
As a final recommendation, the actions for storing the “comparison space” might benefit from following the Memento design pattern [19].
@@ 6.4 Entity Profile Assembly
Because of the large number of ways of aggregating entity references into a coherent entity profile, system designs might benefit from leveraging the Builder design pattern [19] in entity profile assembly.
Certain ER models (like SERF [3]) allow entity profiles to be viewed as entity references. This allows feeding entity profiles back into the system by storing entity profiles is the entity reference store described in Subsection 5.5.
@@ 7.1 Development Considerations
The requirement to implement custom scaling and parallelization mechanisms depends almost entirely on the size of the data being processed. ER systems should leverage existing distributed computing frameworks [40, 36, 60], reducing the complexity of handling large-scale data processing.
@@ 8.1 Concluding Remarks on Research Questions
ER systems must efficiently process diverse input data while maintaining strict output consistency. To support downstream integration, the reference architecture promotes standardization: input flexibility is encouraged, while output representation is constrained. Logical components surround the core ER processes, ensuring uniformity in input reduction via comparison space generation and in entity profile assembly based on ER models. Despite differences in implementation, core entity profile assembly activities remain consistent across ER systems.
@@ 8.3 Final Words
The paper was written in hopes of triggering an effort to standardize ER design across all flavours of ER. Subsequent iterations of this reference architecture (based on feedback and collaboration), and the standardization of the broader ER problem domain would both be welcome.

### v3

# Resolvi: A Reference Architecture for Extensible, Scalable and Interoperable Entity Resolution

@@ Abstract
Context: Entity resolution (ER) plays a pivotal role in data management by determining whether multiple records correspond to the same real-world entity. Because of its critical importance across domains such as healthcare, finance, and machine learning and its long research history designing and implementing ER systems remains challenging in practice due to the wide array of methodologies and tools available. This diversity results in a paradox of choice for practitioners, which is further compounded by the various ER variants (record linkage, entity alignment, merge/purge, a.s.o).
Objective: This paper introduces Resolvi, a reference architecture for facilitating the design of ER systems. The goal is to facilitate creating extensible, interoperable and scalable ER systems and to reduce architectural decision-making duration.
Methods: Software design techniques such as the 4+1 view model or visual communication tools such as UML are used to present the reference architecture in a structured way. Source code analysis and literature review are used to derive the main elements of the reference architecture.
Results: This paper identifies generic requirements and architectural qualities of ER systems. It provides design guidelines, patterns, and recommendations for creating extensible, scalable, and interoperable ER systems. Furthermore, it highlights implementation best practices and deployment strategies based on insights from existing systems.
Conclusion: The proposed reference architecture offers a foundational blueprint for researchers and practitioners in developing extensible, interoperable, and scalable ER systems. Resolvi provides clear abstractions and design recommendations which simplify architecture decision making, whether designing new ER systems or improving existing designs.
journal: Information and Software Technology
@@ 2.1 Related Work
In the absence of a comprehensive review focusing on the disambiguation aspect of named entity recognition and disambiguation, several independent sources were studied to better understand the problem. The opinion that named entity disambiguation systems are designed very similarly to merge/purge systems is rooted in the NERD framework [56] itself, which implements a special ontology-based merge algorithm of extracted named entities. FOX [64] uses building blocks which are prevalent in record linkage systems, too (e.g indexing components, external knowledge bases). Modern systems such as Ensemble Nerd [9] use separate black box subsystems (in this case neural networks) to perform feature extraction and matching, similarly to DeepMatcher [46]. Finally, to the best of our knowledge there is no dispute over the need to extract structured information from unstructured text across the named entity recognition literature.
@@ 2.4 Reference ER Systems
Most of the extensibility built into existing systems focuses on the key aspects of ER (matching, clustering, data preprocessing). The Magellan [30] framework for entity matching is extensible in a completely different way: it allows using matching components such as Ditto [35] or replacing the entire rule-based ER runtime with a full neural network architecture designed for ER such as DeepMatcher [46]. This emphasizes the higher extensibility of ER systems with a framework design.
@@ 5.3 Comparison Space Generator
5.4 Reference Store
The Reference Store appears in both incremental and batch ER, even though it is strictly necessary only for the incremental variant. Existing systems often implement it as an in-memory structure. Common representations include a sequence of pairs of matching entity references (e.g Ditto [35]), a similarity graph (e.g. JedAI [53]) or a logical graph (e.g FAMER [60]). Certain closed-source implementations reportedly persist long-lived identifiers linked to entity references [70], though their exact design remains unknown.
When it is designed as a component, it is reminiscent of the Memento design pattern [19] because it is primarily used to store and access versions of the same data. Common concerns to consider when designing the reference store include data availability and integrity.
@@ 5.5 Matching Engine
Additionally, the matching process can be symbolically reduced to a functional representation, barring the need to inspect the detailed logical components involved. The black box approach to matching was popularized by the Stanford Entity Resolution Framework (SERF) [3]. One critique of the SERF match function is that it implicitly assumes a fixed output structure, whereas in practice, matching strategies differ widely.
@@
System architectures should design the matching engine in the most suitable way to their problem domain and chosen ER model. Certain architectures may benefit from leveraging multiple matching engines. A legitimate concern may regard the ability to scale the matching engine independently of the other components in the ER pipeline because of the higher computational expense inherent in this component. These considerations encourage using the Strategy design pattern [19] in object-oriented designs or functional programming for encapsulating the algorithms which sit at the core of the matching engine. The concerns around scalability and interoperability of the matching engine with its surrounding pipeline components strongly suggest using a component-oriented design.
Finally, the possibility to implement systems that do not leverage similarity, renders the matching engine as an optional component in the ER pipeline.
@@ 5.6 Clustering Engine
Widely regarded as an essential step, not all reference systems implement it transparently. Some ER systems, such as Duke [21] obfuscate clustering by making it part of intermediate storage operations which use its system of MatcherListeners. System developers are informed about clustering possibilities only once the LinkDatabaseMatchListener and LinkDatabase concepts are perused. This obfuscates the pipeline structure of the ER process. On the other hand, some entity matching systems such as Ditto [35] or DeepMatcher [46] focus solely on identifying matching entity references, partially or completely omitting transitive resolution. In some cases, this approach can be beneficial, as it offloads the bad-triplet problem [1] to downstream systems.
At the other end of the spectrum, entity alignment and record linkage systems designed around graph models inherently support clustering. Some ER solutions go further by eliminating the matching phase entirely. For example, clustering subsumes matching by employing functional similarity measures and a carefully chosen merge function (e.g SERF [3]), generic distance metrics [72], or probabilistic models which form clusters directly from the comparison space using similarity functions to model colocation (e.g D-blink [40]).
When applied, clustering typically operates on data supplied by the reference store. In cases where matching is omitted, the entity reference store effectively defines the relevant comparison space. Systems such as the JedAI [53] (entityclustering package) and FAMER [60] (famer-clustering package) already support various clustering algorithms [50, 61], offering flexible options for different ER workflows.
Given its varying adoption across reference systems and the limited research on cluster coherence [70], system architectures may design a clustering engine if their specific problem domain and ER model of choice require one. In object-oriented designs, the Strategy design pattern [19] may be used to encapsulate the various algorithms sitting at the core of the clustering engine. Similarly to the matching engine, component oriented design improves interoperability and, indirectly, scalability.
New system architectures should carefully consider clustering engines for incremental ER. Only some clustering algorithms (such as agglomerative clustering) are generally incremental, i.e can be easily carried out using the batch linkage method [24].
Finally, the evaluation of clustering performance is not as common among the reference ER systems as the fundamental importance of the clustering subsystem to ER would imply. D-blink [40] is an example system which implements this feature by implementing the Adjusted Rand Index [71] for evaluating ER quality. ER system architectures which incorporate clustering engines should also provide the means for evaluating clustering performance along with matching performance, especially since libraries which implement clustering metrics for ER are readily available [49]. Future efforts in this area would improve ER system transparency and reliability.
@@ 6 ER Runtime Process
Up until this point, the paper focused on generic use cases and abstract logical components. Procedurally, ER may be interpreted as a series of subprocesses and activities which, in abstract, remain consistent across existing approaches.
@@ 6.2 Comparison Space Generation
The process of generating the comparison space, denoted as (C) in Figure 10, consists of three main phases: candidate arrangement, comparison input generation, and filtering. These phases largely echo the JedAI Toolkit’s process of blocking, block cleaning and comparison cleaning. Candidate arrangement reorganizes entity references into efficient data structures without loss of information. Techniques such as blocking, indexing, and graph edge modification exemplify this step, resulting in structures like block tables, balanced trees, or graphs. In some cases, these structures may be persistently stored for later retrieval.
@@
As a final remark, the actions for storing the “comparison space” might resemble the Memento design pattern [19].
@@ 6.4 Entity Profile Assembly
Because of the large number of ways of aggregating entity references into a coherent entity profile, entity profile assembly system designs might end up resembling Builder[19].
Certain ER models (like SERF [3]) allow entity profiles to be viewed as entity references. This allows feeding entity profiles back into the system. The reference store (Subsection 5.4) most naturally addresses this need.
@@ 7.1 Development Considerations
The requirement to implement custom scaling and parallelization mechanisms depends almost entirely on the size of the data being processed. ER systems should leverage existing distributed computing frameworks (e.g Dblink [40], Splink [36] or FAMER [60]), reducing the complexity of handling large-scale data processing.
@@ 8.1 Concluding Remarks on Research Questions
ER systems must efficiently process diverse input data while maintaining strict output consistency to support both upstream and downstream integration. Logical components which standardize input and output surround a very flexible and extensible core. This ensures uniformity in input reduction via comparison space generation as well as in entity profile assembly based on ER models. This is possible only because despite differences in implementation, core entity profile assembly activities remain consistent across ER systems.
@@ 8.3 Final Words
The paper was written in hopes of triggering an effort to standardize ER design across all flavours of ER. Both subsequent iterations of this reference architecture (based on feedback and collaboration), and the standardization of the broader ER problem domain would be welcome.